# Università della Svizzera italiana, Switzerland

### v1

@@ 5 Implications of our Findings
Code-LLMs can be safely quantized down to 4-bit precision without compromising performances in source code generation tasks. We empirically showed that quantizing code-LLMs down to 4-bit precision can effectively reduce memory footprint by approximately threefold (x3) without sacrificing performance. This achievement is relevant in different contexts and domains. First, it enables the deployment of larger code models in hardware-constrained environments characterized by limited resources. The consequence is a reduction in energy consumption and an enhancement in terms of scalability, sustainability, and accessibility of advanced AI systems grounded on LLMs for code. Moving along these lines, the use of quantized methods also holds potential in the areas of security and privacy. Outsourcing code to external tools and services to leverage code-LLMs’ automation capabilities can pose significant security risks, including the possibility of data leakage (Novelli et al., 2024; Wu et al., 2024; Huang et al., 2023).
In contrast, reduced models can be run locally thanks to quantization without significantly compromising performance, eliminating the need to depend on third-party solutions and addressing security and privacy concerns. Moreover, for small businesses and startups that lack access to large-scale computing infrastructure, quantization provides an opportunity for innovation and offers a competitive edge, as these organizations can adopt state-of-the-art AI solutions for automating software engineering-related practices (Hou et al., 2024). Another compelling application is in the educational setting, where students may not have access to high-performance computing resources like dedicated servers. A quantized model that can run locally offers a valuable opportunity to enrich education in the software engineering field, giving students access to cutting-edge technology that enhances their learning experience.
Additional research is needed on more extreme quantization levels. While quantization has proven to be an effective tool when bit precision is not pushed to the extreme (i.e., 2-bit), we observed that in cases of ultra-low-level quantization, external factors such as the calibration dataset play a significantly more critical role. Therefore, the influence of the calibration dataset is not negligible and requires careful consideration. Optimizing the structure and composition of the calibration dataset becomes essential in such scenarios, as it can greatly affect the overall performance of the quantized model.
@@
Additionally, with the emergence of very Large Language Models, models with 100 billion parameters or more are becoming the standard rather than the exception, even in software engineering. Codex (Chen et al., 2021), the model behind GitHub Copilot, is based on GPT-3, which boasts over 100 billion parameters. However, the application of quantization to models of this size remains largely unexplored. Investigating quantization at this scale offers significant opportunities and could yield new insights that are not yet fully understood, given the current state of affairs.
6 Threats to Validity
@@
Threats to construct validity concern the relationship between theory and observation. This threat is mainly related to the measurements made to address our research questions. As far as models’ accuracy is concerned, we leverage the pass@k measure, used in previous work (Chen et al., 2021; Cassano et al., 2023, 2024; Wei et al., 2023a) including the Wei et al.study we aimed at partially replicating (Wei et al., 2023a).
@@
Threats to internal validity concern factors internal to our study that could affect our results. These threats mainly concern the choice of hyperparameters for the quantization process, for the end-to-end fine-tuning, and for the inference phase. In Section 3, we explain, for the different phases of our study, how the different hyperparameters have been chosen. More specifically, in some cases, we based our choices on what was done in previous work, including the original AQLM paper (Egiazarian et al., 2024) and the quantization study we replicated (Wei et al., 2023a). In other cases, we made empirical choices, for example when determining the number of epochs in the end-to-end fine-tuning.
@@
Threats to conclusion validity concern the relationship between experimentation and outcome. To support our findings, we leverage suitable statistical procedures, i.e., McNemar’s test (with adjusted p-values using the Benjamini-Hochberg procedure (Yoav and Yosef, 1995)) and Odds Ratio.
@@
Threats to external validity concern the generalizability of our findings. Concerning the studied LLMs, we focused on two widely used models of code, CodeLLama and DeepSeek-Coder. Also, we experimented with different model sizes, up to 34B. However, it is possible that other code-based LLMs or general-purpose LLMs may be differently impacted by the quantization process.
Another threat to external validity concerns the generalizability to other tasks. In this work, we evaluated the effects of quantization on code generation tasks. Therefore, the obtained results may not generalize to other SE-related tasks.
7 Conclusion and Future Work
@@
Previous work by Wei et al. (2023a) showed quantization to be an effective way of managing the size of state-of-the-art LLMs that power AI-driven solutions for the software engineering domain. However, the recent surge in quantization techniques and code-LLMs made us question on what would be the pros and cons of quantization given (i) new advancements, including not only adopting extreme quantization precision levels but also different types of calibration datasets and (ii) large code models featuring up to 34 Billion parameters.
@@
It expands the potential for efficient AI-driven software engineering solutions and paves the way for new research opportunities while providing practical applications for industry professionals.
Our next step is to assess the impact of quantization on a wider range of code-related tasks, such as code summarization and program repair, beyond just code completion. These additional analyses would offer a more comprehensive understanding of how quantization affects code models across the entire spectrum of code-related tasks, thereby helping to generalize our findings further.
Declarations
Fundings
Di Penta acknowledges the Italian “PRIN 2022” project TRex-SE: “Trustworthy Recommenders for Software Engineers”, grant n. 2022LKJWHC. Giagnorio and Bavota thank the Swiss National Science Foundation (SNSF) for the funding provided under the project “PARSED” (grant agreement No. 219294).
Ethical approval
Not applicable.
Informed Consent
Not applicable.
Author Contributions
Alessandro Giagnorio: Developed the research idea, designed the study, conducted experiments, performed statistical analysis, participated in the writing.
Antonio Mastropaolo: Developed the research idea, designed the study, participated in the writing.
Saima Afrin: Developed the research idea, designed the study, participated in the writing.
Massimiliano Di Penta: Developed the research idea, designed the study, participated in the writing.
Gabriele Bavota: Developed the research idea, designed the study, participated in the writing.
Data Availability Statement
Code and data used for conducting the study are publicly available in our replication package (Giagnorio et al., 2025). In particular, we provide (i) the code needed to run quantization, (ii) different quantized models, (iii) all the predictions generated by each quantized model as well as by the baseline, and (iv) additional scripts used in the computation of our findings.
Conflict of Interest
The authors have no conflicts of interest to declare that are relevant to the content of this article.
Clinical trial number
Not applicable.

### v2

@@ 5. Implications of our Findings
Code-LLMs can be safely quantized down to 4-bit precision without compromising performances in source code generation tasks. We empirically showed that quantizing code-LLMs down to 4-bit precision can effectively reduce memory footprint by approximately threefold (x3) without sacrificing performance. This achievement is relevant in different contexts and domains. First, it enables the deployment of larger code models in hardware-constrained environments characterized by limited resources. The consequence is a reduction in energy consumption and an enhancement in terms of scalability, sustainability, and accessibility of advanced AI systems grounded on LLMs for code. Moving along these lines, the use of quantized methods also holds potential in the areas of security and privacy. Outsourcing code to external tools and services to leverage code-LLMs’ automation capabilities can pose significant security risks, including the possibility of data leakage (Novelli et al., 2024; Wu et al., 2024; Huang et al., 2023). In contrast, reduced models can be run locally thanks to quantization without significantly compromising performance, eliminating the need to depend on third-party solutions and addressing security and privacy concerns. Moreover, for small businesses and startups that lack access to large-scale computing infrastructure, quantization provides an opportunity for innovation and offers a competitive edge, as these organizations can adopt state-of-the-art AI solutions for automating software engineering-related practices (Hou et al., 2024). Finally, in the emergent area of agentic AI systems, where multiple smaller models can outperform single monolithic LLMs (Belcak et al., 2025), quantization can effectively enable the deployment of large code models as part of such systems, even in resource-constrained environments.
@@
Additional research is needed on more extreme quantization levels. While quantization has proven to be an effective tool when bit precision is not pushed to the extreme (i.e., 2-bit), we observed that in cases of ultra-low-level quantization, external factors such as the calibration dataset play a significantly more critical role. Therefore, the influence of the calibration dataset is not negligible and requires careful consideration. Optimizing the structure and composition of the calibration dataset becomes essential in such scenarios, as it can greatly affect the overall performance of the quantized model. Additionally, with the emergence of very Large Language Models, models with 100 billion parameters or more are becoming the standard rather than the exception, even in software engineering. Codex (Chen et al., 2021), the model behind GitHub Copilot, is based on GPT-3, which boasts over 100 billion parameters. However, the application of quantization to models of this size remains largely unexplored. Investigating quantization at this scale offers significant opportunities and could yield new insights that are not yet fully understood, given the current state of affairs.
6. Threats to Validity
@@
Construct validity. This threat is mainly related to the measurements made to address our research questions. As far as models’ accuracy is concerned, we leverage the pass@k measure, used in previous work (Chen et al., 2021; Cassano et al., 2023, 2024; Wei et al., 2023) including the Wei et al. study we aimed at partially replicating (Wei et al., 2023).
@@
Internal validity. These threats mainly concern the choice of hyperparameters for the quantization process, for the end-to-end fine-tuning, and for the inference phase. In Section 3, we explain, for the different phases of our study, how the different hyperparameters have been chosen. More specifically, in some cases, we based our choices on what was done in previous work, including the original AQLM paper (Egiazarian et al., 2024) and the quantization study we replicated (Wei et al., 2023). In other cases, we made empirical choices, for example when determining the number of epochs in the end-to-end fine-tuning.
@@
Conclusion validity. To support our findings, we leverage suitable statistical procedures, i.e., McNemar’s test (adjusted p-values via Benjamini-Hochberg procedure (Yoav and Yosef, 1995)) and Odds Ratio.
@@
External validity. Concerning the studied LLMs, we focused on two widely used models of code, CodeLLama and DeepSeek-Coder. Also, we experimented with different model sizes, up to 34B. However, it is possible that other code-based LLMs or general-purpose LLMs may be differently impacted by the quantization process. Although we did not experiment with larger models (e.g., CodeLlama 70B), the trend observed in RQ3 suggests that the benefits of extreme quantization are likely to extend to 34B models as well. Another threat to external validity concerns the generalizability to other tasks. In this work, we evaluated the effects of quantization on code generation tasks. Therefore, the obtained results may not generalize to other SE-related tasks.
7. Conclusion and Future Work
@@
Previous work (Wei et al., 2023) showed quantization to be an effective way of managing the size of state-of-the-art LLMs that power AI-driven solutions for the software engineering domain. However, the recent surge in quantization techniques and code-LLMs made us question on what would be the pros and cons of quantization given (i) new advancements, including not only adopting extreme quantization precision levels but also different types of calibration datasets and (ii) large code models featuring up to 34 Billion parameters.
@@
Our next step is to assess the impact of quantization on a wider range of code-related tasks, such as code summarization and program repair, beyond just code generation. We also plan to investigate the effects of quantization on other dimensions, such as inference latency and energy consumption. These additional analyses would offer a more comprehensive understanding of how quantization affects code models across the entire spectrum of code-related tasks, thereby helping to generalize our findings further.
8. Acknowledgments
Di Penta acknowledges the Italian “PRIN 2022” project TRex-SE: “Trustworthy Recommenders for Software Engineers”, grant n.
2022LKJWHC. Dr. Mastropaolo acknowledges the support of the National Science Foundation (NSF) under grant CCF-2451058.
Giagnorio and Bavota thank the Swiss National Science Foundation (SNSF) for the funding provided under the project “PARSED” (grant agreement No. 219294). Giagnorio also thanks CHOOSE for sponsoring his trip to the conference.